OpenStudy (anonymous):

Use f(x) = 5x − 4 and g(x) = 3 − x^2 to evaluate the expression. (g compose f)(x)

OpenStudy (anonymous):

(g o f )(x) means g(f(x)), the output of f(x) will be the input of g(x) so for the x value of g(x), substitute the function f(x).

OpenStudy (anonymous):

Okay, I don't even know what I'm doing wrong but I'm not coming out to the correct answer according to webassign. I watched their 'watch it' video multiple times and that is not helping me either. 'Watch it' video: Use f(x)=3x-5 and g(x)=2-x^2 to evaluate the expression. (b) (g compose f)(x)=g(f(x)) =g(3x-5) =2-(3x-5)^2 =2-[9x^2-30x+25] =-9x^2+30x-23 I don't even know where they got the 30 from and that is throwing me off.

OpenStudy (akashdeepdeb):

\((a-b)^2 = a^2 - 2ab + b^2\)

OpenStudy (anonymous):

I seriously do not know what I'm doing wrong. But thanks anyways.

OpenStudy (anonymous):

from the video: (g compose f)(x)=g(f(x)) =g(3x-5) =2-(3x-5)^2-> substitute f(x) into g(x) =2-[9x^2-30x+25] -> expand the exponent =-9x^2+30x-23 -> simplify From your problem: 3-(5x-4)^2 -> substitute f(x) into g(x) 3-(25x^2-40x+16) -> expand the exponent 3-25x^2+40x-16 -> simplify -25x^2+40x-13 which part you didn't get??